I. Investigate whether the following functions are even or odd:

(a) f(x) = x3

(b) f(x) = cos x


II. State the mean value theorem

III. (a) Find the derivative of the function y = 2x2 + 12/x2 when x = 2

(b) f(x) = -3/x-7. Find the inverse of the function.

Iv. Consider the function f(x) = erx Determine the values of r so that f satisfies the equation f"(x) + f'(x) - 6f(x) = 0.





1
Expert's answer
2021-04-15T06:50:25-0400
  1. The function "f(x)" is called odd if and only if "-f(x)=f(-x)." It is called even if and only if "f(x)=-f(x)".

(a) The function is odd, since "-x^3=(-x)^3" .

(b) The function is even, since "cos(x)=cos(-x)" .

2. The mean value Theorem states the following: Suppose that function "f(x)" is continuous on the closed interval "[a,b]" and is differentiable on the open interval "(a,b)". Then there exists the point "c\in(a,b)" such that "f'(c)=\frac{f(b)-f(a)}{b-a}".

3. (a) The derivative of the function is: "y'=4x-24x^{-3}". After substituting "x=2" we receive: "y'(2)=8-\frac{24}{8}=8-3=5".

(b) We receive: "f(x)+7=-\frac3x". It yields: "x=-\frac{3}{f(x)+7}". It can be also rewritten as: "f^{-1}(x)=-\frac{3}{x+7}".

4. We substitute "f(x)=e^{rx}" and receive: "r^2+r-6=0" ; "r=\frac{-1\pm5}{2}=2;-3." The solutions are: "f(x)=e^{2x}" and "f(x)=e^{-3x}".